# WateRF: Robust Watermarks in Radiance Fields for Protection of Copyrights

### v1

@@ Abstract
The advances in the Neural Radiance Fields (NeRF) research offer extensive applications in diverse domains, but protecting their copyrights has not yet been researched in depth. Recently, NeRF watermarking has been considered one of the pivotal solutions for safely deploying NeRF-based 3D representations. However, existing methods are designed to apply only to implicit or explicit NeRF representations. In this work, we introduce an innovative watermarking method that can be employed in both representations of NeRF. This is achieved by fine-tuning NeRF to embed binary messages in the rendering process. In detail, we propose utilizing the discrete wavelet transform in the NeRF space for watermarking. Furthermore, we adopt a deferred back-propagation technique and introduce a combination with the patch-wise loss to improve rendering quality and bit accuracy with minimum trade-offs. We evaluate our method in three different aspects: capacity, invisibility, and robustness of the embedded watermarks in the 2D-rendered images. Our method achieves state-of-the-art performance with faster training speed over the compared state-of-the-art methods. https://kuai-lab.github.io/cvpr2024waterf/
@@ 1 Introduction
Digital watermarking plays a pivotal role in reinforcing the copyright of digital assets,  e.g., text, image, audio, video, and 3D content. Digital assets are easily misused without proper admission from the creators and owners of the digital assets. One way to protect the copyright is by encoding invisible watermarks in the digital assets.
Neural Radiance Fields (NeRF) [29] have emerged into the spotlight in 3D content creation and 3D modeling since NeRF can represent 3D objects or scenes in a compact way. With increasing interest in 3D content for the Metaverse, head mount display, virtual reality, and augmented reality, NeRF has become important in digital watermarking research.
@@
A straightforward approach for protecting rendered 3D images from the NeRF model is using the post-generation watermarking method that embeds the watermark into the 2D rendered image from NeRF with the conventional watermarking methods. However, this method only protects copyright for the rendered images but not the copyright of the NeRF model. A fundamental solution to protecting the model and rendered image is fine-tuning the NeRF model to encode the watermark into the NeRF model itself. The watermark will be encoded into the rendered image every time during the rendering process.
NeRF has two primary representations: implicit and explicit representations. The implicit representation [2, 29, 31] uses a multi-layer perceptron (MLP) to represent a 3D scene and the explicit representation [4, 8, 57] uses a traditional 3D data structure, such as voxels. Prior methods [24, 16] were limited to applying to only one type of NeRF representation. However, in order to overcome this limitation, we explore a new question: How can we embed the watermark into two distinct NeRF representations(implicit and explicit)?
In this paper, we propose an innovative watermarking method that can be applied to both types of representations of NeRF. Our method integrates the watermarking process into the rendering process without changing the model architecture. It adjusts the pre-trained NeRF model so that all of the rendered images have a watermark embedded. We pre-train the extractor with the conventional deep learning method HiDDeN [58] to extract the watermark from the rendered image. Our method has several advantages. First, since it is a fine-tuning process, we can deal with the implicit and the explicit NeRF representations. Second, we do not need any extra processes for concealing watermarks in the rendered images. Therefore, less computational cost is required. Third, our method can protect the model and the rendered images simultaneously.
@@
Our extensive experiments demonstrate that our method successfully encodes watermarks by fine-tuning the pre-trained NeRF models and identifying the watermark message using a simplified decoder from HiDDeN [58]. Moreover, we show that our watermarking method is robust from diverse watermark attacks. We evaluate our method by measuring the capacity and robustness under diverse watermark attacks. The training time of our method is approximately six times faster than the CopyRNeRF [24]. Since our method can be applied to both explicit [4] and implicit [29] NeRF, our method can be used in more general applications than the other watermarking methods. Our method outperforms the other state-of-the-art watermarking methods within all metrics. In summary, our contributions are as follows:
@@ 2 Related Work
Digital watermarking. Watermarking methods have been evolving over the decades [39, 3]. There are two main ways to recover watermarks: multi-bit watermark and zero-bit watermark [17]. The multi-bit watermark allows encoding and decoding of multi-bit messages from media. On the other hand, the zero-bit watermark only determines if the media is watermarked. In this paper, we mainly focus on the zero-bit watermark method. More recently, with the great success of deep learning watermark methods like HiDDeN [58], considerable research has been conducted to embed watermarks into carriers especially into 2D images, video and 3D [40, 7, 21, 22, 23, 50].
@@
  • Discrete Fourier Transform (DFT) [48] expresses images in terms of phase and amplitude. When a watermark is embedded into the frequency domain converted by DFT, the magnitude is invariant, making it robust against geometric attacks such as rotation, scaling, and translation [14, 11, 35].
  • Discrete Cosine Transform (DCT) [1, 41] decomposes the energy of image data into a sum of cosine functions by representing signals in the frequency domain. It is useful for image compression because most of the energy of the image is concentrated in the top-left corner coefficient [33].
  • Discrete Wavelet Transform (DWT) [38, 44, 37] decomposes the image into four subbands, LL(Low-Low), LH(Low-High), HL(High-Low), and HH(High-High). The LL band has the most energy and contains low-frequency information. Additionally, the LL band can be recursively decomposed into n-levels. DWT is used in several watermarking methods [49, 6] and shows significant differences in performance depending on the level [9, 47, 15].
In this paper, we use DWT and the LL subband for fine-tuning NeRF to embed the watermark.
Figure 2: Two phases of method. Phase 1: We train the encoder and the decoder to extract messages. After training is done, we do not use the encoder. Phase 2: We fine-tune the NeRF to embed the messages into the rendered images automatically. (a) We disable auto-differentiation and render a full-resolution image in order to save memory. (b) We use DWT for the rendered images and choose the LL coefficient as the input of the pre-trained decoder. (c) We enable auto-differentiation and render the images patch by patch. Then the NeRF is optimized using Eq 7 and Eq 8.
@@ 3 Preliminaries: Representation of NeRF
In this paper, we focus on both implicit and explicit representations of a NeRF model. The implicit NeRF [29] represents a scene using a multi-layer perceptrons $\Phi$ (MLP) whose input is 3D location $\boldsymbol{x}=(x,y,z)\in\mathbb{R}^{3}$ and 2D viewing direction $\boldsymbol{d}\in\mathbb{R}^{2}$ and output is color $\boldsymbol{c}\in\mathbb{R}^{3}$, volume density $\sigma$ $\in\mathbb{R}^{+}$. These MLP-based radiance fields can be written as:
@@
Both representations of NeRF are followed by a volumetric rendering equation, which is used to synthesize novel views:
@@ 4 Method
We propose a process of fine-tuning the NeRF that does not involve alterations to the model’s architecture to embed the watermark message. Our method aims to embed the watermark to the weights $\theta$ of the NeRF model in the frequency domain of rendered images. Our method stands apart from traditional watermarking methods, focusing on training encoders and decoders. The difference lies in the fine-tuning process, which embeds the watermark without using an encoder. There are 2 phases: (1) Pre-training the watermark decoder $D$, (2) Fine-tuning the NeRF model $\mathit{F}_{\theta}$ to embed the message. Our method is illustrated in Fig. 2 and described in detail below.
@@ 4.2 Embedding and extracting watermark on DWT
Watermarking in the spatial domain is a relatively simple method, as it embeds the watermark across the whole image. Recently, a fine-tuning watermarking method for NeRF [16] in the spatial domain rises to the surface. Although the fine-tuning method of embedding a message in the spatial domain shows incomparable invisibility and message extraction ability, it is vulnerable to attacks that distort the spatial domain, such as cropping. Applying spatial domain techniques from the latent diffusion model [7] directly does not allow for effective adjustment of NeRF’s weights.
To tackle these problems, we propose a fine-tuning method in a frequency domain instead of a spatial domain. Various watermarking techniques for images use the frequency domain, including Discrete Cosine Transform (DCT) and Discrete Wavelet Transform (DWT), have seen continuous development and improvement over the years. We find that DWT is the appropriate domain for encoding the message into the weights of a NeRF model.
@@
Using the HiDDeN decoder, we experimentally discover that embedding a watermark in the LL subband is more robust and embeds the watermark information more effectively than other subbands.
The DWT is characterized by its subbands being computed across different levels; therefore, choosing an optimal level for our purpose is necessary. The 1-level separates the images to 4 subbands $(LL_{1},LH_{1},HL_{1},HH_{1})$, then 2-level separate the $LL_{1}$ subband into 4 subbands $(LL_{2},LH_{2},HL_{2},HH_{2})$. We experimentally discover that selecting a level too high decreases visual quality due to excessive adjustment of crucial visual elements. Thus, we use the 2-level DWT because it best maintains a good balance between bit accuracy and reconstruction quality.
@@ 4.3 Fine-tuning the NeRF model
NeRF has two different representations, implicit and explicit representation (see Sec. 3). Prior studies [24, 16] focused on applying their methods to just one form of NeRF representation. In contrast, we propose a method applicable to both NeRF representations: implicit and explicit NeRF representations. Our approach involves fine-tuning the NeRF to ensure that images of the novel views include an embedded message.
@@
Since rendering with NeRF requires a lot of memory, we turn off the auto-differentiation at first. The rendered images $X\in\mathbb{R}^{H\times W\times 3}$ at full resolution are then created by $F_{\theta_{0}}$. Following this, the images $X$ are transformed into seven groups of wavelet subbands $\{LL_{2},LH_{2},HL_{2},HH_{2},LL_{1},LH_{1},HL_{1},HH_{1}\}$. Our pre-trained decoder $D$ takes the $LL_{2}$ subband as an input and extracts a message $M^{\prime}=D(W_{\varphi}(j_{0},m,n))$ from it. The message loss is calculated by comparing $M$ and $M^{\prime}$ using the formula: $L_{m}=BCE(M,sg(M^{\prime}))$. We use the Waston-VGG [5] to calculate the full image perceptual loss between the image with no watermark $X_{0}$ and the image with a watermark $X$: $L_{i}(X_{0},X)$. Full resolution loss function is a sum of the full image perceptual loss and the message loss.
@@
We cache the gradient calculated by Eq. 7 prior to the patch-wise rendering process to update the parameters. Then, we turn on the auto-differentiation so that the NeRF parameters enable patch-wise rendering. Our patch loss $L_{patch}$ is calculated by combining mean absolute error (MAE) across rendered pixel colors, total variation (TV) regularization, and SSIM loss for maintaining the balance between bit-accuracy and visual quality.
@@ 4.4 Deferred back-propagation with patch loss
Our method implement a full-resolution rendering derived from our pre-trained decoder during the fine-tuning process. Each iteration presents a considerable challenge due to the extensive memory consumption. In order to relieve the memory constraints, we adopt a deferred back-propagation [54] specifically designed for memory-efficient, patch-wise optimization within NeRF architectures.
Our method proposes a novel patch loss to be applied to the deferred back-propagation as detailed in Eq. 8. When the model renders the full-resolution image with auto-differentiation disabled, the image and message loss are computed by the rendered image, and the derived gradients are cached (see Sec. 4.3 for more details). Then, the model renders the full-resolution image with image patches that are re-rendered with auto-differentiation enabled and back-propagated using the corresponding gradients from the cached set. Our innovation is not solely relying on cached gradients for back-propagation during re-rendering. Instead, we calculate the loss between the re-rendered patch and the ground truth (GT) patch and carry out an additional back-propagation process. The patch loss consists of a sum of L1 loss (MAE), a Total Variation loss, and an SSIM loss, as defined in Eq. 8. By utilizing double back-propagation, once for the combined loss from the full-resolution image and the message loss and once for the patch loss, we are able to enhance both image quality and message bit accuracy at once.
@@ 5.2 Implementation Details
TensoRF [4] and Vanilla NeRF [29] employ similar training methods, with some specifics as detailed below. The training process is carried out on an image-by-image basis, with a batch size set to one. Ray tracing is implemented along with the camera pose, and the Adam optimizer [12] is used for the learning process. The training is completed with epochs ranging from 5 to 10. Since our pre-trained decoder requires full-resolution images, NeRF renders full-resolution images during the fine-tuning process. However, NeRF requires a great amount of memory usage. We adopt the deferred back-propagation in ARF [54] to solve the memory issue and combine it with the patch-wise loss to improve rendering quality and bit accuracy. The initial $\lambda_{i}$, $\lambda_{m}$ was set to 0.1 and 0.9 for the LLFF dataset and 0.05 and 0.95 for the Blender dataset, respectively. Additionally, within the patch loss equation (Eq. 8), $\lambda_{MAE}$, $\lambda_{TV}$, and $\lambda_{SSIM}$ were set to 0.1, 0.06, and 0.02.
@@ 5.3 Evaluation
We evaluate our method with three key essential aspects in watermarking techniques: 1) invisibility, which involves evaluating the difficulty in detecting embedded watermarks in images by peoples; 2) capacity, measuring the length of the messages that can be embedded into the image; 3) robustness, evaluating the robustness of our watermarking method under various image distortions such as Gaussian Noise, Rotation, Scaling, Gaussian Blur, Crop, Brightness, JPEG compression and combination of these distortions. We employ Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [55] for the evaluation metrics to assess invisibility while utilizing bit accuracy to evaluate capacity and robustness. We conduct experiments on the invisibility and robustness of the message length of $M_{L}\in\{4,8,16,32,48\}$.
Baseline To the best of our knowledge, CopyRNeRF [24] stands out alone in its ability to embed bits with the implicit NeRF model, which can be converted into various modality data such as messages or images. The CopyRNeRF performs state-of-the-art on bit-accuracy and visual quality. As a result, we conduct our evaluations utilizing the CopyRNeRF. We select relevant comparative strategy models as follows: HiDDeN [58]+NeRF [29], which employ the classical 2D watermarking method HiDDeN [58] on training images prior to training the NeRF model.

### v4

@@ Abstract
The advances in the Neural Radiance Fields (NeRF) research offer extensive applications in diverse domains, but protecting their copyrights has not yet been researched in depth. Recently, NeRF watermarking has been considered one of the pivotal solutions for safely deploying NeRF-based 3D representations. However, existing methods are designed to apply only to implicit or explicit NeRF representations. In this work, we introduce an innovative watermarking method that can be employed in both representations of NeRF. This is achieved by fine-tuning NeRF to embed binary messages in the rendering process. In detail, we propose utilizing the discrete wavelet transform in the NeRF space for watermarking. Furthermore, we adopt a deferred back-propagation technique and introduce a combination with the patch-wise loss to improve rendering quality and bit accuracy with minimum trade-offs. We evaluate our method in three different aspects: capacity, invisibility, and robustness of the embedded watermarks in the 2D-rendered images. Our method achieves state-of-the-art performance with faster training speed over the compared state-of-the-art methods. Project page: https://kuai-lab.github.io/cvpr2024waterf/
@@ 1 Introduction
Digital watermarking plays a pivotal role in reinforcing the copyright of digital assets,  e.g., text, image, audio, video, and 3D content. Digital assets are easily misused without permission from the creators and owners of the digital assets. One way to protect the copyright is by encoding invisible watermarks in the digital assets.
Neural Radiance Fields (NeRF) [35] have emerged into the spotlight in 3D content creation and modeling since NeRF can represent 3D objects or scenes in a compact way. With increasing interest in 3D content for the Metaverse, virtual reality, and augmented reality, NeRF has become important in digital watermarking research.
@@
A straightforward approach for protecting rendered 3D images from the NeRF model is using the post-generation watermarking method that embeds the watermark into the 2D rendered image from NeRF with existing watermarking methods. However, this method only protects copyright for the rendered images but not the copyright of the NeRF model. A fundamental solution to protecting the model and rendered image is fine-tuning the NeRF model to encode the watermark into the NeRF model itself. The watermark will be encoded into the rendered image from the fine-tuned NeRF every time during the rendering process.
NeRF has two primary representations: implicit and explicit representations. The implicit representation [1, 35, 37] uses a multi-layer perceptron (MLP) to represent a 3D scene and the explicit representation [3, 10, 58] uses a traditional 3D structure, such as voxels. Prior methods [30, 22] were limited to applying to only one type of NeRF representation. However, in order to overcome this limitation, we explore a new question: How can we embed the watermark into two distinct NeRF representations(implicit and explicit)?
In this paper, we propose an innovative watermarking method that can be applied to both types of NeRF representations. Our method integrates the watermarking process into the rendering process without changing the model architecture. It adjusts the pre-trained NeRF model so that all of the rendered images have same watermark embedded. We pre-train the decoder with the conventional deep learning method HiDDeN [59] to extract the watermark from the rendered image. Our method has several advantages. First, since it is a fine-tuning process, we can deal with the implicit and the explicit NeRF representations. Second, we do not need any extra processes for concealing watermarks in the rendered images. Therefore, less computational cost is required. Third, our method can protect the model and the rendered images simultaneously.
@@
Our extensive experiments demonstrate that our method successfully encodes watermarks by fine-tuning the pre-trained NeRF models and identifying the watermark message using a simplified decoder from HiDDeN [59]. Moreover, we show that our method is robust from diverse watermark attacks. We evaluate our method by measuring the capacity and robustness under diverse watermark attacks. The training time of our method is approximately six times faster than the CopyRNeRF [30]. Since our method can be applied to both implicit [35] and explicit [3] NeRF, our method can be used in more general applications than the other watermarking methods. Our method outperforms the other state-of-the-art watermarking methods within all metrics. In summary, our contributions are as follows:
@@ 2 Related Work
Digital watermarking. Watermarking methods have been evolving over the decades [43, 2]. There are two main ways to recover watermarks [23]: multi-bit watermark and zero-bit watermark. The multi-bit watermark allows for the encoding and decoding of multi-bit messages from media. The zero-bit watermark can be used for ownership authentication. In this paper, we focus on the zero-bit watermark, which involves encoding and decoding multi-bit messages. Recently, with the great success of deep learning watermark methods like HiDDeN [59], considerable research has been conducted to embed watermarks into carriers, especially into 2D images [8, 27, 28], video [29, 7] and 3D [51, 9, 49].
@@
  • Discrete Fourier Transform (DFT) expresses images in terms of phase and amplitude. When a watermark is embedded into the frequency domain converted by DFT, the magnitude is invariant, making it robust against geometric attacks such as rotation and scaling  [20, 14, 41].
  • Discrete Cosine Transform (DCT) decomposes the energy of image data into a sum of cosine functions by representing signals in the frequency domain. It is useful for image compression because most of the energy of the image is concentrated in the top-left corner coefficient. Therefore, DCT is extensively used in watermarking methods that energy compaction. [5, 38, 17]
  • Discrete Wavelet Transform (DWT) decomposes the image into four subbands, LL(Low-Low), LH(Low-High), HL(High-Low), and HH(High-High). The LL band has the most energy and contains low-frequency information. Additionally, the LL band can be recursively decomposed into n-levels. DWT is used in several watermarking methods [50, 6] and shows significant differences in performance depending on the level [11, 48, 21]. In this paper, we use DWT and the LL subband for fine-tuning NeRF to embed the watermark.
Figure 2: WateRF overview. Phase 1: We train the encoder and the decoder to extract messages. After phase 1, we do not use the encoder. Phase 2: We fine-tune the NeRF to embed the messages into the rendered images. (a) We disable auto-differentiation and render a full-resolution image to save memory. (b) We use DWT for the rendered images and choose the LL subband as the input of the pre-trained decoder. (c) We enable auto-differentiation and render the images patch by patch. Then the NeRF is optimized using Eq 7 and Eq 8.
@@ 3 Preliminaries: Representation of NeRF
In this paper, we focus on both implicit and explicit representations of a NeRF model. The implicit NeRF [35] represents a scene using a multi-layer perceptrons (MLP) $\Phi$ whose input is 3D location $\boldsymbol{x}=(x,y,z)\in\mathbb{R}^{3}$ and 2D viewing direction $\boldsymbol{d}\in\mathbb{R}^{2}$ and output is color $\boldsymbol{c}\in\mathbb{R}^{3}$, volume density $\sigma$ $\in\mathbb{R}^{+}$. These MLP-based radiance fields can be written as:
@@
Both representations are followed by a volumetric rendering equation, which is used to synthesize novel views:
@@ 4 Method
We propose a process of fine-tuning the NeRF that does not involve alterations to the model’s architecture to embed the watermark message. Our method aims to embed the watermark into the weights $\theta$ of the NeRF model in the frequency domain of rendered images. Our method stands apart from traditional watermarking methods, which focusing on training encoders and decoders. The difference lies in the fine-tuning process, which embeds the watermark without using an encoder. There are two phases: (1) Pre-training the watermark decoder $D$, (2) Fine-tuning the NeRF model $\mathit{F}_{\theta}$ to embed the message. Our method is illustrated in Fig. 2 and described in detail below.
@@ 4.2 Embedding and extracting watermark on DWT
Recently, a fine-tuning watermarking method for NeRF [22] in the spatial domain has emerged. Although the fine-tuning method of embedding a message in the spatial domain shows incomparable invisibility and message extraction ability, it is vulnerable to attacks that distort the spatial domain, such as cropping. Directly applying spatial domain techniques from the latent diffusion model [8] does not allow for effective adjustment of NeRF’s weights.
To tackle these problems, we propose a fine-tuning method in a frequency domain instead of a spatial domain. Various watermarking techniques for images use the frequency domain have seen development and improvement over the years. We find that DWT is the appropriate domain for encoding the message into the weights of a NeRF model.
@@
Previous studies [15, 18, 13] have selected $LH$, $HL$, and $HH$ subbands for embedding watermarks because the $LL$ subband contains significant information about the image. However, we choose the LL subband as an input of our decoder D and get the extract message with $M^{\prime}=D(W_{\varphi})$. We experimentally discover that embedding the watermark in the LL subband is more robust and effective than other subbands for the HiDDeN decoder.
‘The DWT is characterized by its subbands being computed across different levels; therefore, choosing an optimal level for our purpose is necessary. The 1-level separates the images to 4 subbands $(LL_{1},LH_{1},HL_{1},HH_{1})$, then 2-level separate the $LL_{1}$ subband into 4 subbands $(LL_{2},LH_{2},HL_{2},HH_{2})$. We experimentally discover that selecting a level too high decreases visual quality due to excessive adjustment of crucial visual elements. Thus, we use the 2-level DWT because it best maintains a good balance between bit accuracy and reconstruction quality.
@@ 4.3 Fine-tuning the NeRF model
NeRF has two different representations, implicit and explicit representations (see Sec. 3). Previous studies [30, 22] focused on applying their methods to just one form of NeRF representation. In contrast, we propose a method applicable to both NeRF representations: implicit and explicit. Our approach involves fine-tuning the NeRF to ensure that novel view images include an embedded message.
@@
Since rendering with NeRF requires a lot of memory, we turn off the auto-differentiation at first. The rendered images $X\in\mathbb{R}^{H\times W\times 3}$ at full resolution are rendered by $F_{\theta_{0}}$. Following this, the full resolution images $X$ are transformed into seven groups of wavelet subbands : $\{LL_{2},LH_{2},HL_{2},HH_{2},LL_{1},LH_{1},HL_{1},HH_{1}\}$. Our pre-trained decoder $D$ takes the $LL_{2}$ subband as an input and extracts a message $M^{\prime}=D(W_{\varphi}(j_{0},m,n))$ from it. The message loss is calculated by comparing $M$ and $M^{\prime}$: $L_{m}=BCE(M,sg(M^{\prime}))$. We choose the Waston-VGG [4] to calculate the full image perceptual loss between the image without watermark $X_{0}$ and the image with a watermark $X$: $L_{i}(X_{0},X)$. Full resolution loss function is a sum of the full image perceptual loss and the message loss.
@@
We cache the gradient calculated by Eq. 7 prior to the patch-wise rendering process to update the parameters. Then, we turn on the auto-differentiation so that the NeRF parameters enable patch-wise rendering. Our patch loss $L_{patch}$ is calculated by combining mean absolute error (MAE) across rendered pixel colors, total variation (TV) regularization, and SSIM loss to maintain the balance between bit-accuracy and visual quality.
@@ 4.4 Deferred back-propagation with patch loss
Our method implements a full-resolution rendering derived from our pre-trained decoder during the fine-tuning process. Each iteration presents a considerable challenge due to the extensive memory consumption. In order to relieve the memory constraints, we adopt the deferred back-propagation [55], specifically designed for memory-efficient, patch-wise optimization within NeRF.
Our method proposes a novel patch loss to be applied to the deferred back-propagation as detailed in Eq. 8. When the model renders the full-resolution image with auto-differentiation disabled, the image and message loss are computed by the rendered image, and the derived gradients are cached (see Sec. 4.3 for more details). Then, the model renders the full-resolution image with image patches that are re-rendered with auto-differentiation enabled and back-propagated using the corresponding gradients from the cached set.
Our innovation is not solely relying on cached gradients for back-propagation during re-rendering. Instead, we calculate the loss between the re-rendered patch and the ground truth (GT) patch and carry out an additional back-propagation process. The patch loss consists of a sum of L1 loss (MAE), a Total Variation loss, and an SSIM loss, as defined in Eq. 8. By utilizing back-propagation, once for the combined loss from the full-resolution image and the message loss and once for the patch loss, we are able to enhance both image quality and message bit accuracy at once.
@@ 5.2 Implementation Details
Vanilla NeRF [35] and TensoRF [3] employ similar training methods, with some specifics as detailed below. The training process is carried out on an image-by-image basis, with a batch size set to one. Ray tracing is implemented along with the camera pose, and the Adam optimizer [16] is used for the learning process. The training is completed with epochs ranging from 5 to 10. Since our pre-trained decoder requires full-resolution images, NeRF renders full-resolution images during the fine-tuning process. However, NeRF requires a great amount of memory usage. We adopt the deferred back-propagation in ARF [55] to solve the memory issue and combine it with the patch-wise loss to improve rendering quality and bit accuracy.
The initial $\lambda_{i}$, $\lambda_{m}$ was set to 0.1 and 0.9 for the LLFF dataset and 0.05 and 0.95 for the Blender dataset, respectively. Additionally, within the patch loss equation (Eq. 8), the parameters $\lambda_{MAE}$, $\lambda_{TV}$, and $\lambda_{SSIM}$ were set to 0.1, 0.06, and 0.02.
@@ 5.3 Evaluation
We evaluate our method with three key essential aspects in watermarking techniques: 1) capacity, measuring the length of the messages that can be embedded into the image; 2) invisibility, which involves evaluating the difficulty in detecting embedded watermarks in images by peoples; 3) robustness, evaluating the robustness of our watermarking method under various image distortions such as Gaussian Noise, Rotation, Scaling, Gaussian Blur, Crop, Brightness, JPEG compression and combination of these distortions. We employ Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [56] for the evaluation metrics to assess invisibility while utilizing bit accuracy to evaluate capacity and robustness. We conduct experiments on the invisibility and robustness of the message length of $M_{L}\in\{4,8,16,32,48\}$.
Baseline To the best of our knowledge, CopyRNeRF [30] stands out alone in its ability to embed bits with the implicit NeRF model, which can be converted into various modality data such as messages or images. The CopyRNeRF performs state-of-the-art on bit-accuracy and visual quality. As a result, we conduct our evaluations utilizing the CopyRNeRF. We select the following relevant comparative models:: HiDDeN [59]+NeRF [35], which employ the classical 2D watermarking method HiDDeN [59] on training images prior to training the NeRF model.